A copper washer is to be fit in place over a steel
Answer #1

Givan

thermal expansion for copper and steel is 16x10^-6/K and 12x10^-6/K respectively

Expansion = coefficient of expansion x temperature change

the washer needs to expand 0.0030 cm. (1.3660 - 1.3630)

0.0030 = 0.000016 * ( T(final) - 23.2)

T(final) - 23.2 = 0.0030 / 0.000016

T(final) - 23.2= 187.5

T(final) = 187.5 + 23.2

T(final) = 210.7 degrees C